OpenStudy (anonymous):

Is the dot product: vector (dot) vector= scalar? cross product = vector (dot) vector = vector?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

well, the dot in cross should be an x

OpenStudy (anonymous):

cross product is for both 2-d and 3-d?

OpenStudy (anonymous):

@amistre64 oh yes,

OpenStudy (amistre64):

\[dot:\vec u \cdot \vec v=n\]\[cross:\vec u \times \vec v=\vec s\]

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

can dot product be scalar (dot) scalar? or scalar(dot) vector?

OpenStudy (amistre64):

the operations are defined on vectors ... you can relate it to a linear vector space (1 0) if need be

OpenStudy (amistre64):

hmmm, by applying the definition of a dot product we can produce the elements needed for the resultant vector \[\binom{1}{0}(-3)=\binom{-3\cdot 1}{-3\cdot 0}=\binom{-3}{0}\] but the operations are not defined particularly for non vector spaces

OpenStudy (amistre64):

an R^1 vectors looks like a number, yes but it has to be viewed in the context of its 1-dimentional vector space

OpenStudy (amistre64):

\[\vec u=(3)~:~\vec v=(-4)\] \[\vec u\cdot \vec v = 3(-4)=-12\] \[\vec u\times \vec v = 0\]

OpenStudy (anonymous):

how did it become zero?

OpenStudy (amistre64):

\[det\begin{pmatrix}3&-4\\0&0\end{pmatrix}=3(0)+4(0)=0\]

OpenStudy (amistre64):

or rather \[\vec 0\]

OpenStudy (anonymous):

oh..I see thanks

OpenStudy (amistre64):

another way to consider it is that in R^1, vectors either in the same direction, or 180 degrees opposites in either case, they are scalar multiples of each other and have a zero determinant in the cross

OpenStudy (anonymous):

thank you...

OpenStudy (amistre64):

youre welcome